Jane is the sole shareholder of Buttons, Inc. Buttons has accumulated earnings and profits (E & P) of $65,000 at the beginning of the current year. The current E & P is $35,000. Buttons pays out a property distribution to Jane during the current year with an FMV of $150,000 and an adjusted basis of $130,000. How much is taxable dividend to Jane?

a. $35,000

b. $100,000

c. $120,000

d. $150,000

Please show your work.
Answer #1

Option c. 120,000 is correct.

Distribution to the extent of balance in E&P is taxable as dividend.

Gain recognized by Buttons on distribution increases E&P. Gain = 150,000 - 130,000 = 20,000

Total E&P balance = Accumulated E&P + Current E&P + gain on distribution = 65,000 + 35,000 + 20,000 = 120,000.

Thus, taxable dividend is 120,000 (as distribution of $150,000 is higher than E&P balance)